OpenStudy (anonymous):

Helps guys please fast! 3. graph the function and identify the vertical asymptote. y=1/x-4-2 4. graph the function and identify the vertical asymptote. y= 1/x+2+4

OpenStudy (anonymous):

are you a quick learner or you just want answers

OpenStudy (anonymous):

both work just fine

OpenStudy (anonymous):

how can both work?

OpenStudy (anonymous):

im a quick learner.but u can give me the answers to

OpenStudy (anonymous):

i see so where are you stuck

OpenStudy (anonymous):

i dont know how to solve it..;?

OpenStudy (anonymous):

put values of x and y as infinity to get them respectively

OpenStudy (anonymous):

i kinda need the answers fast tho

OpenStudy (anonymous):

because a curve tends towards its asymptotes as it approaches infinity

OpenStudy (anonymous):

lol !

OpenStudy (anonymous):

@aviz SO WOULD the answer be x=-4

OpenStudy (anonymous):

for the first one/

OpenStudy (anonymous):

first (y+2) = 1/(x-4) or (x-4)=1/(y+2), as y tends to infinity x-4=0, x=4

OpenStudy (anonymous):

can u help with the second one to please?! and thank you

OpenStudy (anonymous):

@aviz

OpenStudy (anonymous):

each one has two asymptotes x=4 and y=-2 and second one x=-2 and y=4

OpenStudy (anonymous):

vertical ones are the x ones